OpenStudy (trojanpoem):

Where is the ICR ? I found two velocities IC = 4/6 = 2/3 ft (1-2/3) * 6 = 2 ft/s (1+2/3) * 6 = 10 ft/s which is A and which is B Yes it's slipping.

OpenStudy (trojanpoem):

Solve

OpenStudy (irishboy123):

This is Galilean Relativity Freeze the conveyor, the disc is then moving to the left at 16 in that ref frame. And still rotating at 6. then compare 16 and 6* 1

OpenStudy (trojanpoem):

You wanna solve it using Galilean transformation ?

OpenStudy (mathmate):

 At IC, the instantaneous velocity =6* (2/3)-4=0, hence IC. For A, V=6*1-4=2 For B, V=-6*1-4=-10

OpenStudy (trojanpoem):

Why IC under B not above b ?

OpenStudy (mathmate):

Correction on labeling.